# Model Predictive Control for Aggressive Driving Over Uneven Terrain

### v2

@@ I Introduction
First, rollover can occur when the vehicle becomes disabled by tipping onto its sides. The risk of rollovers is substantially higher in off-road driving where steep hills, evasive sharp turns, and vehicles with high centers of mass due to tall suspension systems are common. The second failure risk is associated with ditches. Often characterized by extreme impact forces, snapped or irreparably bent control arms (fig. 1.2) are common failures due to poor ditch traversal. Paper Website: https://sites.google.com/cs.washington.edu/off-road-mpc
@@ II Related Work
Currently, there are few works which conduct autonomous off-road experimentation on a full-scale platform, fewer which demonstrate speeds beyond 7 $\mathrm{m}\mathrm{/}\mathrm{s}$, and, to the authors’ best knowledge, none which additionally deploy in real, uneven terrain. However, work on the same chassis, sensor, and compute platform does exist [16, 15, 45]. These works focus primarily on individual sub-systems, validate with offline data and methods, and do not report on system performance or speeds. Our work presents the most complete description and evaluation of a control system for the fastest full-scale, off-road, autonomous vehicle in uneven terrain to date. For more distant but related work, we note that due to lack of standardization in the off-road setting, difficulty of the environment and expected performance can vary drastically. Prior works on real, full-scale vehicles do not attempt to navigate in uneven terrain [7, 37]. However, even in mild terrain conditions, our operating speeds of 7–10 $\mathrm{m}\mathrm{/}\mathrm{s}$ are higher or comparable to other work. In [44], the system navigates autonomously at about 3.2 $\mathrm{m}\mathrm{/}\mathrm{s}$ and averages 4 interventions over 1.6 km. In the DARPA Grand Challenge, the fastest method achieves speeds between 9–11 $\mathrm{m}\mathrm{/}\mathrm{s}$ on largely flat terrain, slowing to 4-6 $\mathrm{m}\mathrm{/}\mathrm{s}$ for mild bumps [37]. Older work also demonstrate capabilities up to 9.7 $\mathrm{m}\mathrm{/}\mathrm{s}$ over flat ground and binary traversability classification [7].
@@
Existing work on real systems on uneven terrain but at low speeds generally assume the vehicle adheres to the ground at all times and constrain attitude angles as estimated by only the ground geometry [9, 11, 24]. This method is relatively fast, simple to implement, and effective at low speeds [19]. However, at higher speeds, the method becomes increasingly inaccurate and unsafe due to dynamic effects from unmodeled forces and vehicle inertia. On the other hand, complex dynamics models can explictly predict wheel forces and trajectories but are often too slow for safe operation. For instance, [50] implements a 3D rigid-body model with tire forces in MPC at a control rate of 3 Hz, about one tenth of our target rate. In [39], the control rate is sufficient at 50 Hz but operates at speeds of 3 $\mathrm{m}\mathrm{/}\mathrm{s}$ over gentler and differentiable ground manifolds. Note that due to the architecture of many off-road perception systems, ground manifolds are often outputs of complex neural networks [32, 15] or approximations to point clouds [24]. Neither representation admits a practical path toward real-time optimization using aforementioned methods.
Some works have pursued learning-based approaches to driving through uneven terrain. In [4], the authors propose using imitation learning for safe traversal of various types of ditches at 1 $\mathrm{m}\mathrm{/}\mathrm{s}$ in simulation. However, success rates drop by as much as 50% depending on what types of train and test geometries are selected. Due to the continuum of terrain geometry in natural environments, categorizing ditches and assuming dataset coverage may be impractical and unsafe. Learning-based dynamics modeling has also been proposed in the literature [16, 43, 49]. However, in the aggressive regime, collecting data near the safety limits of the vehicle can be dangerous and infeasible in practice. In [30], the authors seek to generalize low speed training data to higher speed tests using physics-based modeling, improving tight turns on flat ground over a basic bicycle model.
@@ III-B Rollover
A final vector triple product identity $(\mathbf{a}\times\mathbf{b})\times\mathbf{c}=(\mathbf{a}\cdot\mathbf{c})\mathbf{b}-(\mathbf{a}\cdot\mathbf{b})\mathbf{c}$ allows further simplification,
@@ V-C2 Course Time
The course time is the total time it takes the vehicle to complete the course by reaching every waypoint in order—minus the time used to reset the vehicle due to any interventions. An aim of aggressive driving is to (safely) reduce travelling time.
@@ V-C3 Ditch False-Positives
A falsely identified ditch (false-positive or FP) is a common complication in the development of a ditch control method due in part to imperfect perception. Because safe ditch traversal speeds are usually around 1-3 $\mathrm{m}\mathrm{/}\mathrm{s}$, unnecessary slowdowns from our nominal speeds of 7-10 $\mathrm{m}\mathrm{/}\mathrm{s}$ are detrimental to aggressive autonomy.
We find that the geometry-based baseline is more susceptible to ditch FPs despite best efforts in parameter tuning to maintain safe ditch traversal while reducing FPs. It is likely due to the fact that purely geometrical reasoning treats all detected ditches equally regardless of current speed or actual severity. Moreover, there is less flexibility in the tunable parameter space to compensate for this, since we have shown in section III the severity is necessarily correlated with traversal speeds. The accumulation of multiple FPs inevitably increase the course time on the Tour for the baseline. On the Ruot, our method encounters a potential FP while ascending a steep hill and stalling due to slip before the crest. This consequently results in an intervention so as not to damage the transmission belt. It is possible the behavior was due to a FP which would slow the vehicle down at the crest. Without slippage modeling, the predicted velocity would have resulted in potential airtime and violation of (12). As the Baja is mostly ditch traversals and obstacle avoidance, it is unlikley to experience a false-positive during the course due to the density of ditches. However, some of the ditches are not so severe as to require full slowdowns. The baseline method does not accommodate this and thus considerably increases the completion time due to over-compensation of these ditches.
@@ Acknowledgment
This research was developed with funding from the Defense Advanced Research Projects Agency (DARPA). The views, opinions and/or findings expressed are those of the author and should not be interpreted as representing the official views or policies of the Department of Defense or the U.S. Government.
@@ VII-B3 Costmap Costs
TABLE III: Table of relevant parameter settings. The temperature, number of samples, and horizon are with respect to MPPI as in [48].

### v3

@@ I Introduction
First, rollover can occur when the vehicle becomes disabled by tipping onto its sides. The risk of rollovers is substantially higher in off-road driving where steep hills, evasive sharp turns, and vehicles with high centers of mass due to tall suspension systems are common. The second failure risk is associated with ditches. Often characterized by extreme impact forces, snapped or irreparably bent control arms (fig. 1.2) are common failures due to poor ditch traversal.
@@ II Related Work
Currently, there are few works which conduct autonomous off-road experimentation on a full-scale platform, fewer which demonstrate speeds beyond 7 $\mathrm{m}\mathrm{/}\mathrm{s}$, and, to the authors’ best knowledge, none which additionally deploy in real, uneven terrain. However, work on the same chassis, sensor, and compute platform does exist [16, 15, 45]. These works focus primarily on individual sub-systems, validate with offline data and methods, and do not report on system performance or speeds. Our work presents the most complete description and evaluation of a control system for the fastest full-scale, off-road, autonomous vehicle in uneven terrain to date. For more distant but related work, we note that due to lack of standardization in the off-road setting, difficulty of the environment and expected performance can vary drastically. Prior works on real, full-scale vehicles do not attempt to navigate in uneven terrain [7, 37]. However, even in mild terrain conditions, our operating speeds of 7–10 $\mathrm{m}\mathrm{/}\mathrm{s}$ are higher or comparable to other work. In [44], the system navigates autonomously at about 3.2 $\mathrm{m}\mathrm{/}\mathrm{s}$ and averages 4 interventions over 1.6 km. In the DARPA Grand Challenge, the fastest method achieves speeds between 9–11 $\mathrm{m}\mathrm{/}\mathrm{s}$ on largely flat terrain, slowing to 4-6 $\mathrm{m}\mathrm{/}\mathrm{s}$ for mild bumps [37]. Older work also demonstrate capabilities up to 9.7 $\mathrm{m}\mathrm{/}\mathrm{s}$ over flat ground and binary traversability [7].
@@
Existing work on real systems on uneven terrain but at low speeds generally assume the vehicle adheres to the ground at all times and constrain attitude angles as estimated by only the ground geometry [9, 11, 24]. This method is relatively fast, simple to implement, and effective at low speeds [19]. However, at higher speeds, the method becomes increasingly inaccurate and unsafe due to dynamic effects from unmodeled forces and vehicle inertia. On the other hand, complex dynamics models can explictly predict wheel forces and trajectories but are often too slow for safe operation. For instance, [50] implements a 3D rigid-body model with tire forces in MPC at a control rate of 3 Hz, about one tenth of our target rate. In [39], the control rate is sufficient at 50 Hz but operates at speeds of 3 $\mathrm{m}\mathrm{/}\mathrm{s}$ over gentler and differentiable ground manifolds. Note that due to the architecture of common off-road perception systems, elevation maps are often outputs of complex neural networks [32, 15] or approximations to point clouds [24]. Neither representation admits a practical path toward real-time optimization using aforementioned methods.
Some works have pursued learning-based approaches to driving through uneven terrain. In [4], the authors propose using imitation learning for safe traversal of various types of ditches at 1 $\mathrm{m}\mathrm{/}\mathrm{s}$ in simulation. However, success rates drop by as much as 50% depending on which types of train and test geometries are selected. Learning-based dynamics modeling has also been proposed in the literature [16, 43, 49]. However, in the aggressive regime, collecting data near the safety limits of the vehicle can be dangerous and infeasible in practice. In [30], the authors seek to generalize low speed training data to higher speed tests using physics-based modeling, improving tight turns on flat ground over a basic bicycle model.
@@ III-B Rollover
A final vector triple product identity $(\mathbf{a}\times\mathbf{b})\times\mathbf{c}=(\mathbf{a}\cdot\mathbf{c})\mathbf{b}-(\mathbf{a}\cdot\mathbf{b})\mathbf{c}$ allows further algebraic simplification,
@@ V-C2 Course Time
The course time is the total time it takes the vehicle to complete the course by reaching every waypoint in order. An aim of aggressive driving is to (safely) reduce travelling time.
@@ V-C3 Ditch False-Positives
A falsely identified ditch (false-positive or FP) is a common complication in the development of a ditch control method due in part to imperfect perception. Because safe ditch traversal speeds are typically around 1-3 $\mathrm{m}\mathrm{/}\mathrm{s}$, unnecessary slowdowns from our nominal speeds of 7-10 $\mathrm{m}\mathrm{/}\mathrm{s}$ are detrimental to aggressive autonomy.
We find that the geometry-based baseline is more susceptible to ditch FPs despite best efforts in parameter tuning to maintain safe ditch traversal while reducing FPs. It is likely due to pure geometrical reasoning, which treats all detected ditches equally regardless of current speed or actual severity. Moreover, there is less flexibility in the tunable parameter space to compensate for this, since we have shown in section III the severity is necessarily correlated with traversal speeds. The accumulation of multiple FPs inevitably increases the course time on the Tour for the baseline. With our method on the Ruot, we flag a potential FP encountered when cresting a steep hill (waypoint 5 to 4). Without slippage modeling, MPPI believes its command speed will result in airtime (12) at the crest. In reality, the vehicle lacks momentum and traction to cleanly crest the hill and begins to stall. An intervention is made to prevent damage to the transmission belt. On the Baja, which involves primarily ditches and obstacle avoidance, false-positives are less likely due to the dense presence of ditches. However, some of these ditches do not require full slowdowns for safe traversal. The baseline method does not accommodate this and results in longer course completion times due to over-slowing.
@@ Acknowledgment
This research was developed with funding from the Defense Advanced Research Projects Agency (DARPA). The views, opinions and/or findings expressed are those of the author and should not be interpreted as representing the official views or policies of the Department of Defense or the U.S. Government. This material is also partially based upon work supported by the National Science Foundation Graduate Research Fellowship under Grant No. DGE-2140004. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation.
@@ VII-B3 Costmap Costs
TABLE III: Table of relevant parameter settings. The temperature, number of samples, and horizon are with respect to the MPPI algorithm as laid out in [48].